At temperature T Kelvin (K), the value of the Fermi function at an energy 0.5 eV above the Fermi energy is 0.01. Then T, to the nearest integer, is_____________
(
1
Expert's answer
2019-04-18T11:30:43-0400

The Fermi function f(E) gives the probability that a given available electron energy state will be occupied at a given temperature.

The Fermi function f(E) is given by formula


"f(E)= \frac {1} {1+e^{\frac {E-E_F}{kT}}} (1)"

where EF is the Fermi energy or Fermi level, k is the Boltzmann constant (k=8.6 10-5 eV/K), T is the absolute temperature in K

In our case, f(E)= 0.01 eV, E-EF=0.5 eV

Using (1) we got:


"1+e^{\frac {E-E_F}{kT}}=100"

"T=\frac {0.5}{8.6\times 10^{-5}}=5814 K"

Answer:

5814 K